OpenStudy (anonymous):

show all work in dividing (y^2+5y-6/y^2-3y-10) by (y^2+2y-3/y^2-4y-5)

OpenStudy (anonymous):

\[\frac{ y^2+5y-6 }{ y^2-3y-10 } \div \frac{ y^2+2y-3 }{ y^2-4y-5 }\]

OpenStudy (anonymous):

Hint:\[\frac{ y^2+5y-6 }{ y^2-3y-10 } \div \frac{ y^2+2y-3 }{ y^2-4y-5 }=\frac{(y+6)(y-1)}{(y-5)(y-2)} \div \frac{ (y+3)(y-1)}{ (y-5)(y+1) }\]